OpenStudy (anonymous):

Solve: −5(x − 2) = 15 1 −3.4 −5 −1

OpenStudy (anonymous):

-5(x-2) = 15 -5x +10 =15 -5x = 15-10 -5x = 5 x = 5/-5 x= -1